# Collective Superradiance: Estimating the Peak Emission Rate and Time

### v2

@@ I Introduction
Quantum emitter ensembles interacting through a shared electromagnetic reservoir can exhibit cooperative behavior in the form of superradiance [1, 2, 3, 4, 5, 6, 7]. The case of initially fully inverted ensembles results in the cooperative enhancement of the radiative decay process and the arrival of a superradiant emission peak (see Fig. 1(b)). Accurately predicting the peak photon emission rate and time of such systems is not only a long-standing open problem in theoretical quantum optics but relevant in a wide range of research areas including superradiant lasing [8, 9, 10, 11], biochemistry [12], photochemistry [13, 14], radio astronomy [15, 16], nanomaterials [17], condensed matter physics [18], and quantum error correction [19, 20]. Superradiance in the few- and many-excitation regime has been observed numerous times over the years in a broad range of platforms such as atomic gases [21, 22, 23, 24, 25, 26], solid-state systems [27, 28, 29, 30, 31, 32], molecular emitters [33, 34], ensembles of nuclei [35] and waveguide platforms [27, 36], illustrated in Fig. 1(a). However, theoretically predicting the peak radiated power and time presents a significant computational challenge: tracking the time evolution of the density operator for an ensemble of $N$ quantum emitters, each undergoing collective spontaneous decay, quickly becomes intractable as $N$ increases due to the exponential growth in system dimension. The difficulty lies in the need to account for nonlinear all-to-all interactions between the two-level emitters, inherent in the superradiant decay process. Conventional methods that rely on direct simulation scale exponentially with $N$ [37, 38, 39], often requiring supercomputing resources for even moderately sized emitter numbers. In the simplest case of Dicke’s original superradiance formulation [1], all emitters are assumed to be indistinguishable, allowing even for analytic solutions [40, 41]. For superradiance in extended ensembles [42, 43] numerical methods either rely on approximations in terms of emitter correlations [44, 45, 46], or on symmetries of the problem to reduce complexity [47], however generally the system’s dimension scales polynomially with the emitter number at best. Furthermore, variations in the system’s physical parameters, such as geometry, emitter spacing, and emitter types, add layers of complexity to the problem.
@@
The ability to provide accurate and instant predictions of peak emission properties in large-scale superradiant quantum systems, irrespective of the system configuration, significantly accelerates research and insights into cooperative light-matter platforms ranging from microscopic to macroscopic emitter numbers.
Figure 1: (a) Collective superradiance and correlated decay can occur in scenarios such as in free-space atomic emitter ensembles, molecular emitter ensembles with additional non-collective decay, and quantum emitters coupled to waveguide reservoirs. (b) The emission rate $R(t)$ exhibits a peak value $R_{\mathrm{peak}}$ at $t_{\mathrm{peak}}>0$ for collectively decaying emitters: The result is shown at the example of a ring of atomic emitters in free-space with nearest-neighbor spacing $a=0.1\lambda_{0}$, a chain of $N=10$ molecular emitters with $\Gamma^{\prime}=2\Gamma$, $a=0.1\lambda_{0}$, quantum emitters coupled to a one-dimensional waveguide with $\Gamma^{\prime}=2\Gamma$, relative phase between emitters $k_{\mathrm{0}}a=1.04\pi$ and the Dicke superradiance limit, $a=0$. $N=10$ in all curves. (c) Normalizing the peak emission rate by $\Gamma gN^{g}$, where $g\equiv g^{(2)}(0)$ is the second-order correlation function in Eq. 3, unveils the universal connection between $g$ and $R_{\mathrm{peak}}$. The peak emission rates can be described by a single curve $p(g)$, described in Section III and Appendix A. The data points in (c) are obtained from quantum master equation simulations by varying $a$, $N$, and $\Gamma^{\prime}$ with emitter ensembles in 1D and 2D. The peak emission rates in (b) are indicated by vertical arrows in (c).
@@ II Theoretical description
Table 1: A summary of possible decay processes in an ensemble of $N$ two-level quantum emitters and their associated values of the second-order correlation function $g^{(2)}(0)$ in Eq. 3, or equivalently, the variance of the eigenvalues of the collective dissipation matrix $\mathbf{\Gamma}$.
In the limiting case of a fully excited but independent emitter ensemble, where the emitter separation (far) exceeds $\lambda_{0}$, the peak emission rate occurs at $t=0$ with $R_{\mathrm{peak}}=N\Gamma$ and is followed by an exponential decay. In the limit of small emitter separations below $\lambda$ the correlations in Eq. 2 become non-zero and positive, leading to a rapid release of photons and in the formation of a superradiant burst with $R_{\mathrm{peak}}>N\Gamma$ at a later time $t>0$.
Whether a superradiant peak will appear at $t>0$ has, until recently, been an open question. Recent works however derived a simple condition based either on the second-order correlation function at zero time-delay being larger than one [42, 54, 55] or the time-derivative of the emission rate in Eq. 2 being positive [43], both evaluated at $t=0$. In both approaches the evaluation can be reduced to finding the variance of the eigenvalues of the dissipation matrix $\mathbf{\Gamma}$ presented below.
@@
where $\{\Gamma_{k}\}$ are the eigenvalues of the real symmetric dissipation matrix $\mathbf{\Gamma}$. The previous works were able to obtain precise conditions for superradiance based on the values of Eq. 3 and Eq. 4, summarized in Table 1.
However, obtaining the actual values of the peak emission rate and the time still involved solving the early time evolution of $\rho(t)$ either using the quantum master equation, whose computational cost scales exponentially, or approximate methods which scale at best polynomially with the number of emitters. In the next section, we present two compact formulas, for accurately predicting the peak emission rate and time, only involving the evaluation of the variance in Eq. 4.
Figure 2: Dicke superradiance. Comparison of Eq. 5 and Eq. 6 (dots) with the exact numerics based on the quantum master equation (continuous lines) as a function of $N$. (a) The exact value of the peak emission rate shows excellent agreement with Eq. 5, where $g=2-2/N$. The peak emission rate converges to $\Gamma N^{2}/5$ for large $N$. (b) The peak emission time, and also shown is the time $\tau_{d}=\ln N/(\Gamma N)$ [2] (gray dashed line).
@@ III Superradiance Process
We assume that the system is in initially in the product state $|\psi_{0}\rangle=\bigotimes_{n=1}^{N}|e_{n}\rangle$ with its time evolution governed by Eq. II. If the decay process is superradiant (see Table 1), the emission peak will occur at a delayed time $t>0$, as opposed to exponential or weakly correlated decay, where the maximum emission occurs at $t=0$. To find the peak emission rate and time means tracking the dissipative time dynamics and solving the superradiant decay problem with the emitter density operator $\rho$.
The key result of this work is the following observation: The peak emission rates $R_{\mathrm{peak}}$ shown in Fig. 2(b) depend strongly on the microscopic details of the emitter ensemble, such as geometry, density, type of emitter, dipole orientation. However, upon normalization by $gN^{g}$, we find in Fig. 2(c) that the peak emission rates of any emitter ensemble configuration and emitter type follow the same curve as a function of the second-order correlation function $g^{(2)}(0)$, in Eq. 3. By introducing the short-hand notation $g\equiv g^{(2)}(0)$, the estimation for the peak emission rate reads
@@
Figure 3: Superradiance peak emission rate as a function of emitter separation or number calculated via the quantum master equation (dots) and Eq. 5 (continuous lines). (a) Free-space emitters in a ring configuration, with $a=0.1\lambda_{0}$ on the left and in a $3\times 4$ array on the right. (b) Molecular emitters with the same geometry as in (a) with $a=0.1\lambda_{0}$ on the left and non-collective rates $\Gamma^{\prime}=\Gamma$ (blue) and $\Gamma^{\prime}=2\Gamma$ (orange). (c) Emitter chain coupled to a single-mode waveguide with $k_{0}a=0.02\pi$ on the left and $N=12$ on the right as well as decay rates $\Gamma^{\prime}=2\Gamma$ (blue) and $\Gamma^{\prime}=4\Gamma$ (orange) into non-guided modes. In plots (a) and (b) the emitters are circularly polarized in the array plane. The quantum master equation simulations in Eq. II inlcude the Hamiltonian, but in Eq. 5 it is not accounted for. The influence of the Hamiltonian is treated in Appendix D.
@@
Figure 4: Time of peak emission as a function of emitter separation or number calculated via the quantum master equation (dots) and Eq. 6 (continuous lines) for the same parameters as in Fig. 3. (a) Free-space emitter ring, (b) Molecular emitters and (c) Emitter chain coupled to a single-mode waveguide reservoir. The upper gray dashed line indicates the time $\tau_{d}=\ln N/(N\Gamma)$ [2, 48].
Dicke superradiance – The paradigmatic example studied by Dicke [1] of $N$ indistinguishable two-level emitters, leads to $\Gamma_{nm}=\Gamma$ for all $n,m$ and furthermore we assume $\Gamma^{\prime}=0$. Under these conditions, Eq. 3 simplifies to $g=2-2/N$ and using Eq. 5, the peak emission rate for $N\geq 8$ is approximately given by
@@
using $p(2-2/N)\approx 11/10-g/2$ for $g\gtrsim 1.75$ (see Appendix A) and converges to $\Gamma N^{2}/5$ for large $N$, shown in Fig. 2(a). Eq. 4 results in $\mathrm{Var}({\{\Gamma_{k}\}/\Gamma})=N-1$ and thus in a peak emission time $\Gamma\tau_{\mathrm{peak}}=\ln(N-1)/(N+1)$, close to the delay time $\ln N/(\Gamma N)$ presented in Ref. [2]. Both show excellent agreement in Fig. 2(b) with exact numerics for the large $N$ limit, while the expression obtained here also agrees well at small $N$. We remark, that for emitters with non-collective decay, Dicke superradiance occurs only for $N-2>\Gamma^{\prime}/\Gamma$ [55].
@@
Figure 5: Universal length scaling and bound for free-space superradiance in (a) chain, (b) square and (c) cubic ensembles of two-level emitters. The variance in Eq. 4 determines the exponent $g$ for superradiant peak emission in Eq. 5 and exhibits a universal dependence on the length scale $L=\bar{a}N^{1/\mathcal{D}}$, where $\mathcal{D}$ is the ensemble dimension and $\bar{a}$ is the average emitter spacing. The average spacing is obtained by initializing perfect lattices with lattice spacings $a$ and randomly disorder emitter positions according to a normal distribution of standard deviation $\sigma/a=2$ (see Appendix C). All plots are shown as a function of the spacing $a/\lambda_{0}=[10^{-4},10^{1}]$ and in (d) as a function of the emitter number $N_{y}=1,...,20$. The gray regions indicate the ensemble lengths with $\mathrm{Var(\{\Gamma_{k}\}/\Gamma)/N}\leq 10^{-3}$, where the superradiant peak emission reduces to $R_{\mathrm{peak}}\approx N\Gamma$. (d) For elongated ensembles we find maximal superradiant enhancement for a side length of $L_{y}\approx 0.2\lambda_{0}$ irrespective of emitter number or spacing. Dipole orientations are $\mathbf{d}=(0,0,1)$ in all plots, however, the same conclusions and qualitatively similar curves are obtained considering other dipole orientations in the limit of large $N$. Bounds on the ensemble lengths and expressions for the variances $\mathrm{Var}_{\mathrm{1D,2D,3D}}$ are shown in Table 2.
@@
Emitter ensembles in free-space – Quantum emitter ensembles in free-space decay collectively and superradiantly when their average nearest-neighbor separation is small enough, where the minimal spacing generally depends on the dimensionality of the ensemble [42, 54]. Here we assume identical quantum emitters with $\Gamma^{\prime}=0$ in various ordered configurations, however, the results apply equally to disordered ensembles, presented below and in Appendix C. In Fig. 3(a) peak emission rates are shown for ring, chain, and 2D square geometries, as a function of emitter number and spacing. The exact numerics (dots) based on the quantum master equation in Eq. II show good agreement with Eq. 5. At small spacings, $a\lesssim 0.05\lambda_{0}$ deviations appear, due to the influence of the Hamiltonian in the decay dynamics (see Appendix D). The corresponding peak emission times are plotted in Fig. 4(a) with the same parameters.
@@
Molecular and solid-state emitters – Molecular or solid-state emitters in the excited electronic state decay both collectively ($\Gamma$) via the zero-phonon line (ZPL) and a manifold of vibrational, rotational or phononic states with a total rate quantified by $\Gamma^{\prime}$ and illustrated in Fig. 1(a) [33]. This leads to a quantum efficiency (QE) of the emitter $\mathrm{QE}={\Gamma}/({\Gamma+\Gamma^{\prime}})$ which reduces Eq. 3 and thus leads to a reduced superradiant emission (Fig. 1(b)). However, this is counteracted by the ability to realize deeply subwavelength ensembles with molecular aggregates reaching spacing on the order of tens of nanometers [13, 14], creating the condition for large numbers of emitters to decay collectively. Fig. 3(b) and Fig. 4(b) show comparisons of peak emission rates and times between exact numerics for up to $N=12$ emitter with the quantum master equation. The non-collective rates are $\Gamma^{\prime}=1\Gamma$ and $\Gamma^{\prime}=2\Gamma$, corresponding to quantum efficiencies $QE=50\%$ and $QE\approx 33\%$ respectively. Similar values can be found for instance in Dibenzoterrylene (DBT) molecules embedded in organic nanocrystals [33].
@@
Single-mode waveguide reservoir – Two-level quantum systems coupled to a single-mode (bidirectional) waveguide are illustrated in Fig. 1(a). The coupling efficiency $\beta$ is defined as the ratio of the radiative decay rate of an individual emitter into the waveguide mode $\Gamma$ to its total decay rate $\Gamma+\Gamma^{\prime}$, namely $\mathcal{\beta}={\Gamma}/({\Gamma+\Gamma^{\prime}})$ [53]. The system consists of a chain of emitters positioned along the waveguide at positions $\{x_{n}\}$. For clarity, we assume an ordered chain with nearest-neighbor separation $a$, while in Appendix C we present the case of ensembles with positional disorder. The variance in Eq. 4 now simplifies to [55]
@@
Table 2: Linear chain, square, and cubic ensembles in free-space assuming an equal emitter number in each dimension and with average nearest-neighbor spacing $\bar{a}$. The variance can be approximated by a simple expression depending only on the average nearest-neighbor spacing $\bar{a}$ and total emitter number $N$. Here $\lceil x\rceil$ is the ceiling function and we add the condition $\lceil x\rceil=N$ if $x>N$ such that it is valid for all spacings. In Fig. 5 (a)-(c) we find excellent agreement with the exact numerical value of the variance. We define the maximum ensemble length $L$ as the value where $\mathrm{Var(\{\Gamma_{k}\}/\Gamma)/N}\sim 10^{-3}$, at which point $R_{\mathrm{peak}}\approx N\Gamma$ according to Eq. 3 and Eq. 4 (shown in Fig. 5). This sets an upper limit on the spatial extent of emitter ensembles with 3D dipole-dipole interactions, such as in free space.
Note, that the variance depends only on two parameters, the emitter number $N$ and spacing $a$, allowing us to estimate the peak emission rate and time for large $N$. In Fig. 3(c) and Fig. 4(c) we plot the comparison between exact numerics and the formulas in Eq. 5 and Eq. 6 for linear chains of emitters. Superradiant decay emerges in particular if the relative phase is close to integer multiples of $\pi$, the Dicke limit, as seen by the appearance of a finite delay time in Fig. 4(c). Lastly, as opposed to its free-space counterpart, the variance has a minimum value of $N/2-1$, namely the emitters always decay superradiantly if the condition $N/2>1+\Gamma^{\prime}/\Gamma$ is fulfilled.
We note, that the Hamiltonian $\mathcal{H}$ is not accounted for in Eq. 5 and Eq. 6 and generally leads to a reduction of peak emission at spacings $a<0.1\lambda_{0}$, however its influence decreases with emitter number as shown in the Supplement D.

### v4

# Analytical Prediction of the Superradiant Emission Peak and Timing in Fully Excited Quantum Emitter Ensembles

@@ I Introduction
Quantum emitter ensembles interacting through a shared electromagnetic reservoir can exhibit cooperative behavior in the form of superradiance [1, 2, 3, 4, 5, 6, 7]. The case of initially fully inverted ensembles results in the cooperative enhancement of the radiative decay process and the arrival of a superradiant emission peak (see Fig. 1(b)). Accurately predicting the peak photon emission rate and time of such systems is not only a long-standing open problem in theoretical quantum optics but relevant in a wide range of research areas including superradiant lasing [8, 9, 10, 11], biochemistry [12], photochemistry [13, 14], radio astronomy [15, 16], nanomaterials [17], condensed matter physics [18], and quantum error correction [19, 20]. Superradiance in the few- and many-excitation regime has been observed numerous times over the years in a broad range of platforms such as atomic gases [21, 22, 23, 24, 25, 26], solid-state systems [27, 28, 29, 30, 31, 32], molecular emitters [33, 34], ensembles of nuclei [35] and waveguide platforms [27, 36], illustrated in Fig. 1(a). However, theoretically predicting the peak radiated power and time presents a significant computational challenge: tracking the time evolution of the density operator for an ensemble of $N$ quantum emitters, each undergoing collective spontaneous decay, quickly becomes intractable as $N$ increases due to the exponential growth in system dimension. The difficulty lies in the need to account for nonlinear all-to-all interactions between the two-level emitters, inherent in the superradiant decay process. Conventional methods that rely on direct simulation scale exponentially with $N$ [37, 38, 39], often requiring supercomputing resources for even moderately sized emitter numbers. In the simplest case of Dicke’s original superradiance formulation [1], all emitters are assumed to be indistinguishable, allowing even for analytic solutions [40, 41]. For superradiance in extended ensembles [42, 43] numerical methods either rely on approximations in terms of emitter correlations [44, 45], or on symmetries of the problem to reduce complexity [46], however generally the system’s dimension scales polynomially with the emitter number at best. Furthermore, variations in the system’s physical parameters, such as geometry, emitter spacing, and emitter types, add layers of complexity to the problem.
@@
The ability to provide accurate and instant predictions of peak emission properties in large-scale superradiant quantum systems, irrespective of the system configuration [48], significantly accelerates research and insights into cooperative light-matter platforms ranging from microscopic to macroscopic emitter numbers. For a fully excited initial state, the authors show in Ref. [xxx] that for atomic emitters coupled to three-dimensional photonic environments (i.e. free-space arrays), the peak emission rate scales linearly with emitter number for any configuration outside the Dicke limit.
Figure 1: (a) Collective superradiance and correlated decay can occur in scenarios such as in free-space atomic emitter ensembles, molecular emitter ensembles with additional non-collective decay, and quantum emitters coupled to waveguide reservoirs. (b) The emission rate $R(t)$ exhibits a peak value $R_{\mathrm{peak}}$ at $t_{\mathrm{peak}}>0$ for collectively decaying emitters: The result is shown at the example of a ring of atomic emitters in free-space with nearest-neighbor spacing $a=0.1\lambda_{0}$, a chain of $N=10$ molecular emitters with $\Gamma^{\prime}=2\Gamma$, $a=0.1\lambda_{0}$, quantum emitters coupled to a one-dimensional waveguide with $\Gamma^{\prime}=2\Gamma$, relative phase between emitters $k_{\mathrm{0}}a=1.04\pi$ and the Dicke superradiance limit, $a=0$. $N=10$ in all curves. (c) Normalizing the peak emission rate by $\Gamma gN^{g}$, where $g\equiv g^{(2)}(0)$ is the second-order correlation function in Eq. (3), unveils the universal connection between $g$ and $R_{\mathrm{peak}}$. The peak emission rates can be described by a single curve $p(g)$, described in Section III and Appendix A. The data points in (c) are obtained from quantum master equation simulations by varying $a$, $N$, and $\Gamma^{\prime}$ with emitter ensembles in 1D and 2D. The peak emission rates in (b) are indicated by vertical arrows in (c).
@@ II Theoretical description
Table 1: A summary of possible decay processes in an ensemble of $N$ two-level quantum emitters and their associated values of the second-order correlation function $g^{(2)}(0)$ in Eq. (3), or equivalently, the variance of the eigenvalues of the collective dissipation matrix $\mathbf{\Gamma}$.
In the limiting case of a fully excited but independent emitter ensemble, where the emitter separation (far) exceeds $\lambda_{0}$, the peak emission rate occurs at $t=0$ with $R_{\mathrm{peak}}=N\Gamma$ and is followed by an exponential decay. In the limit of small emitter separations below $\lambda$ the correlations in Eq. (2) become non-zero and positive, leading to a rapid release of photons and in the formation of a superradiant burst with $R_{\mathrm{peak}}>N\Gamma$ at a later time $t>0$.
Whether a superradiant peak will appear at $t>0$ has, until recently, been an open question. Recent works however derived a simple condition based either on the second-order correlation function at zero time-delay being larger than one [42, 54, 55] or the time-derivative of the emission rate in Eq. (2) being positive [43], both evaluated at $t=0$. In both approaches the evaluation can be reduced to finding the variance of the eigenvalues of the dissipation matrix $\mathbf{\Gamma}$ presented below.
@@
where $\{\Gamma_{k}\}$ are the eigenvalues of the real symmetric dissipation matrix $\mathbf{\Gamma}$. The previous works were able to obtain precise conditions for superradiance based on the values of Eq. (3) and Eq. (4), summarized in Table 1.
However, obtaining the actual values of the peak emission rate and the time still involved solving the early time evolution of $\rho(t)$ either using the quantum master equation, whose computational cost scales exponentially, or approximate methods which scale at best polynomially with the number of emitters. In the next section, we present two compact formulas, for accurately predicting the peak emission rate and time, only involving the evaluation of the variance in Eq. (4).
Figure 2: Dicke superradiance. Comparison of Eq. (5) and Eq. (6) (dots) with the exact numerics based on the quantum master equation (continuous lines) as a function of $N$. (a) The exact value of the peak emission rate shows excellent agreement with Eq. (5), where $g=2-2/N$. The peak emission rate converges to $\Gamma N^{2}/5$ for large $N$. (b) The peak emission time, and also shown is the time $\tau_{d}=\ln N/(\Gamma N)$ [2] (gray dashed line).
@@ III Superradiance Process
We assume that the system is in initially in the product state $|\psi_{0}\rangle=\bigotimes_{n=1}^{N}|e_{n}\rangle$ with its time evolution governed by Eq. (II). If the decay process is superradiant (see Table 1), the emission peak will occur at a delayed time $t>0$, as opposed to exponential or weakly correlated decay, where the maximum emission occurs at $t=0$. To find the peak emission rate and time means tracking the dissipative time dynamics and solving the superradiant decay problem with the emitter density operator $\rho$.
The key result of this work is the following observation: The peak emission rates $R_{\mathrm{peak}}$ shown in Fig. 2(b) depend strongly on the microscopic details of the emitter ensemble, such as geometry, density, type of emitter, dipole orientation. However, upon normalization by $gN^{g}$, we find in Fig. 2(c) that the peak emission rates of any emitter ensemble configuration and emitter type follow the same curve as a function of the second-order correlation function $g^{(2)}(0)$, in Eq. (3). By introducing the short-hand notation $g\equiv g^{(2)}(0)$, the estimation for the peak emission rate reads
@@
Figure 3: Superradiance peak emission rate as a function of emitter separation or number calculated via the quantum master equation (dots) and Eq. (5) (continuous lines). (a) Free-space emitters in a ring configuration, with $a=0.1\lambda_{0}$ on the left and in a $3\times 4$ array on the right. (b) Molecular emitters with the same geometry as in (a) with $a=0.1\lambda_{0}$ on the left and non-collective rates $\Gamma^{\prime}=\Gamma$ (blue) and $\Gamma^{\prime}=2\Gamma$ (orange). (c) Emitter chain coupled to a single-mode waveguide with $k_{0}a=0.02\pi$ on the left and $N=12$ on the right as well as decay rates $\Gamma^{\prime}=2\Gamma$ (blue) and $\Gamma^{\prime}=4\Gamma$ (orange) into non-guided modes. In plots (a) and (b) the emitters are circularly polarized in the array plane. The quantum master equation simulations in Eq. (II) inlcude the Hamiltonian, but in Eq. (5) it is not accounted for. The influence of the Hamiltonian is treated in Appendix D.
@@
Figure 4: Time of peak emission as a function of emitter separation or number calculated via the quantum master equation (dots) and Eq. (6) (continuous lines) for the same parameters as in Fig. 3. (a) Free-space emitter ring, (b) Molecular emitters and (c) Emitter chain coupled to a single-mode waveguide reservoir. The upper gray dashed line indicates the time $\tau_{d}=\ln N/(N\Gamma)$ [2, 47].
Dicke superradiance – The paradigmatic example studied by Dicke [1] of $N$ indistinguishable two-level emitters, leads to $\Gamma_{nm}=\Gamma$ for all $n,m$ and furthermore we assume $\Gamma^{\prime}=0$. Under these conditions, Eq. (3) simplifies to $g=2-2/N$ and using Eq. (5), the peak emission rate for $N\geq 8$ is approximately given by
@@
using $p(2-2/N)\approx 11/10-g/2$ for $g\gtrsim 1.75$ (see Appendix A) and converges to $\Gamma N^{2}/5$ for large $N$, shown in Fig. 2(a). Eq. (4) results in $\mathrm{Var}({\{\Gamma_{k}\}/\Gamma})=N-1$ and thus in a peak emission time $\Gamma\tau_{\mathrm{peak}}=\ln(N-1)/(N+1)$, close to the delay time $\ln N/(\Gamma N)$ presented in Ref. [2]. Both show excellent agreement in Fig. 2(b) with exact numerics for the large $N$ limit, while the expression obtained here also agrees well at small $N$. We remark, that for emitters with non-collective decay, Dicke superradiance occurs only for $N-2>\Gamma^{\prime}/\Gamma$ [55].
@@
Figure 5: Universal length scaling and bound for free-space superradiance in (a) chain, (b) square and (c) cubic ensembles of two-level emitters. The variance in Eq. (4) determines the exponent $g$ for superradiant peak emission in Eq. (5) and exhibits a universal dependence on the length scale $L=\bar{a}N^{1/\mathcal{D}}$, where $\mathcal{D}$ is the ensemble dimension and $\bar{a}$ is the average emitter spacing. The average spacing is obtained by initializing perfect lattices with lattice spacings $a$ and randomly disorder emitter positions according to a normal distribution of standard deviation $\sigma/a=2$ (see Appendix C). All plots are shown as a function of the spacing $a/\lambda_{0}=[10^{-4},10^{1}]$ and in (d) as a function of the emitter number $N_{y}=1,...,20$. The gray regions indicate the ensemble lengths with $\mathrm{Var(\{\Gamma_{k}\}/\Gamma)/N}\leq 10^{-3}$, where the superradiant peak emission reduces to $R_{\mathrm{peak}}\approx N\Gamma$. (d) For elongated ensembles we find maximal superradiant enhancement for a side length of $L_{y}\approx 0.2\lambda_{0}$ irrespective of emitter number or spacing. Dipole orientations are $\mathbf{d}=(0,0,1)$ in all plots, however, the same conclusions and qualitatively similar curves are obtained considering other dipole orientations in the limit of large $N$. Bounds on the ensemble lengths and expressions for the variances $\mathrm{Var}_{\mathrm{1D,2D,3D}}$ are shown in Table 2.
@@
Emitter ensembles in free-space – Quantum emitter ensembles in free-space decay collectively and superradiantly when their average nearest-neighbor separation is small enough, where the minimal spacing generally depends on the dimensionality of the ensemble [42, 54]. Here we assume identical quantum emitters with $\Gamma^{\prime}=0$ in various ordered configurations, however, the results apply equally to disordered ensembles, presented below and in Appendix C. In Fig. 3(a) peak emission rates are shown for ring, chain, and 2D square geometries, as a function of emitter number and spacing. The exact numerics (dots) based on the quantum master equation in Eq. (II) show good agreement with Eq. (5). At small spacings, $a\lesssim 0.05\lambda_{0}$ deviations appear, due to the influence of the Hamiltonian in the decay dynamics (see Appendix D). The corresponding peak emission times are plotted in Fig. 4(a) with the same parameters.
@@
Molecular and solid-state emitters – Molecular or solid-state emitters in the excited electronic state decay both collectively ($\Gamma$) via the zero-phonon line (ZPL) and a manifold of vibrational, rotational or phononic states with a total rate quantified by $\Gamma^{\prime}$ and illustrated in Fig. 1(a) [33]. This leads to a quantum efficiency (QE) of the emitter $\mathrm{QE}={\Gamma}/({\Gamma+\Gamma^{\prime}})$ which reduces Eq. (3) and thus leads to a reduced superradiant emission (Fig. 1(b)). However, this is counteracted by the ability to realize deeply subwavelength ensembles with molecular aggregates reaching spacing on the order of tens of nanometers [13, 14], creating the condition for large numbers of emitters to decay collectively. Fig. 3(b) and Fig. 4(b) show comparisons of peak emission rates and times between exact numerics for up to $N=12$ emitter with the quantum master equation. The non-collective rates are $\Gamma^{\prime}=1\Gamma$ and $\Gamma^{\prime}=2\Gamma$, corresponding to quantum efficiencies $QE=50\%$ and $QE\approx 33\%$ respectively. Similar values can be found for instance in Dibenzoterrylene (DBT) molecules embedded in organic nanocrystals [33].
@@
Single-mode waveguide reservoir – Two-level quantum systems coupled to a single-mode (bidirectional) waveguide are illustrated in Fig. 1(a). The coupling efficiency $\beta$ is defined as the ratio of the radiative decay rate of an individual emitter into the waveguide mode $\Gamma$ to its total decay rate $\Gamma+\Gamma^{\prime}$, namely $\mathcal{\beta}={\Gamma}/({\Gamma+\Gamma^{\prime}})$ [53]. The system consists of a chain of emitters positioned along the waveguide at positions $\{x_{n}\}$. For clarity, we assume an ordered chain with nearest-neighbor separation $a$, while in Appendix C we present the case of ensembles with positional disorder. The variance in Eq. (4) now simplifies to [55]
@@
Table 2: Linear chain, square, and cubic ensembles in free-space assuming an equal emitter number in each dimension and with average nearest-neighbor spacing $\bar{a}$. The variance can be approximated by a simple expression depending only on the average nearest-neighbor spacing $\bar{a}$ and total emitter number $N$. Here $\lceil x\rceil$ is the ceiling function and we add the condition $\lceil x\rceil=N$ if $x>N$ such that it is valid for all spacings. In Fig. 5 (a)-(c) we find excellent agreement with the exact numerical value of the variance. We define the maximum ensemble length $L$ as the value where $\mathrm{Var(\{\Gamma_{k}\}/\Gamma)/N}\sim 10^{-3}$, at which point $R_{\mathrm{peak}}\approx N\Gamma$ according to Eq. (3) and Eq. (4) (shown in Fig. 5). This sets an upper limit on the spatial extent of emitter ensembles with 3D dipole-dipole interactions, such as in free space.
Note, that the variance depends only on two parameters, the emitter number $N$ and spacing $a$, allowing us to estimate the peak emission rate and time for large $N$. In Fig. 3(c) and Fig. 4(c) we plot the comparison between exact numerics and the formulas in Eq. (5) and Eq. (6) for linear chains of emitters. Superradiant decay emerges in particular if the relative phase is close to integer multiples of $\pi$, the Dicke limit, as seen by the appearance of a finite delay time in Fig. 4(c). Lastly, as opposed to its free-space counterpart, the variance has a minimum value of $N/2-1$, namely the emitters always decay superradiantly if the condition $N/2>1+\Gamma^{\prime}/\Gamma$ is fulfilled.
We note, that the Hamiltonian $\mathcal{H}$ is not accounted for in Eq. (5) and Eq. (6) and generally leads to a reduction of peak emission at spacings $a<0.1\lambda_{0}$, however its influence decreases with emitter number as shown in the Supplement D.